OpenStudy (anonymous):

x^2+5x-14=0 quadratic equations

OpenStudy (unklerhaukus):

A quadratic Equation\[ax^2+bx+c=0\] The Quadratic formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] Factored form of the quadratic equation\[(x-x_1)(x-x_2)=0\]

OpenStudy (unklerhaukus):

start by identifying the coefficients \(a,b,c\) Compare \[x^2+5x-14=0\]With\[ax^2+bx+c=0\] \[a=?\quad b=?\quad c=?\]

OpenStudy (anonymous):

Thank you

OpenStudy (unklerhaukus):

what are a, b, c @tobi274 ?